(D3 − 3DD′2 + 2D′3)z = 0
1
Expert's answer
2020-06-01T18:35:35-0400

 The auxillary equation is


"m^3-3m+2=0"

"m^3-m-2m+2=0"

"m(m^2-1)-2(m-1)=0"

"(m-1)(m^2+m-2)=0"

"(m-1)^2(m+2)=0"

The roots of the auxiliary equation are "1,1,-2."

The solution is given by


"z=f(y+x)+xg(y+x)+h(y-2x)"